5. a) Define viscosity.Β 

b) A matter has a body-centered cubic unit cell with a length of 330 π‘π‘š along an edge

and a density of 8.57 𝑔 π‘π‘šβˆ’3

. Calculate the atomic mass of the matter.


c) Find the total energy of a moving object of mass 6.3 π‘˜π‘” and momentum

11.5 Γ— 106 π‘˜π‘”π‘šπ‘ 

βˆ’1

.


1
Expert's answer
2021-12-12T16:40:23-0500

a) Viscosity is a property of fluid to resist the changes in its density.

b) The atomic mass:


"M=\frac{\rho N_aa^3}{2}=0.093\text{ kg/mol}."

c) The energy is


"E=\frac{p^2}{2m}=1.2·10^{20}\text{ J}."